A unicycle wheel with a diameter of 0.400 [m] starts rotating from rest with constant angular acceleration 2.00 [rad/s^2]. What is the tangential speed of the wheel’s rim after two revolutions?
1
Expert's answer
2020-12-17T07:26:37-0500

"v=\omega r=\sqrt{2 \epsilon \theta}r=\sqrt{2 \cdot 2 \cdot 4 \pi}\cdot 0.4=2.84 ~ \frac ms."